# Nonstabilizerness Estimation using Graph Neural Networks

### v1

@@ 4.1 Data for Stabilizer State Classification
Note that, since PS is balanced between the two labels and because nonstabilizerness is invariant under Clifford, CS and ES are also balanced. Furthermore, by construction, the circuits in CS preserve part of the structure of the circuits in PS, while the circuits in ES may result in strongly different quantum circuit structures due to the random placement of CNOT gates rather than their sequential appending. The computation of $M_{2}$ for circuits in PS can be easily performed by leveraging its additivity and compute $M_{2}$ as the sum of the $M_{2}$ computed per each qubit. Although CS inherits the $M_{2}$ values from the PS subset, the ES subset does not, since the contribution of each gate to the overall $M_{2}$ can change after the random CNOT injections, in the general case.
We generate two datasets, both including all three circuit type. The first is with a fixed number of qubit ($n=18$), which is generated with the same size of the dataset used in [mello2025retrieving], in order to have a fair comparison between our GNN and CNN. The second dataset instead enables more general experiments including quantum circuits with number of qubits ranging from $2$ to $25$. A schematic summary of the dataset for stabilizer state classification is provided in Table 1.
@@
The presence of CS circuits in the dataset allows to investigate the generalization capabilities of the models to unseen entangled states, while keeping the same labels and $M_{2}$ values, in line with the procedure used in [mello2025retrieving]. The Entangled States (ES) subset, on the other hand, enables a further step in this analysis by testing model performance on circuits characterized by substantially different structural patterns and, in general, a higher degree of entanglement. Figure 1 shows the heterogeneous quantum circuit dataset generated with our procedure, spanning a broad range of qubit number, circuit depth, circuit structure, entanglement and $M_{2}$. Figure 1(a) illustrates the diversity of quantum circuits across the three different types in terms of gate counts and $M_{2}$. It is important to note that our procedure generates also quantum circuits with significantly low $M_{2}$ values. For instance, in the case of $18$-qubit circuits, the minimum $M_{2}$ value obtained in our dataset is $0.01$, compared to $1.80$ reported in [mello2025retrieving]. Moreover, in the dataset PS 2-25, we observe a minimum $M_{2}$ value of $0.006$, see Figure 1(b). This is relevant, as the performance of machine learning models on the stabilizer state classification problem is typically correlated with the density of $M_{2}$ in the number of qubits $n$: $m_{2}=\frac{M_{2}}{n}$. Models tend to achieve higher accuracy for states with higher $m_{2}$, because of the increased separability of the magic states from the stabilizer states. This effect is also shown in the experimental results discussed in [mello2025retrieving] and further confirmed by our own results obtained using graph neural networks (GNNs).
@@
4.2 Data for SRE-based Classification
All stabilizer states are removed from the original dataset, leaving only the quantum circuits labeled as magic states (with label one). Subsequently, the median of the $M_{2}$ values of the remaining circuits is computed and taken as the threshold $M_{2}$ to distinguish between circuits with low and high $M_{2}$. This procedure makes the classification task particularly challenging, as the median corresponds to a region densely populated by quantum circuits with similar $M_{2}$ values. The violin plots in Figure 2 illustrates the distribution of $M_{2}$ values across the product-state datasets. The horizontal lines in the plot represent the median $M_{2}$ value for each dataset shown along the x-axis. As expected, quantum circuits with qubit numbers between $2$ and $10$ exhibit a narrower range of $M_{2}$ values compared to circuits with qubit numbers between $11$ and $25$.
@@ 4.2 Data for SRE-based Classification
We note that by choosing the median as the threshold $M_{2}$ value in our dataset we ensure that is balanced between the two labels: low-$M_{2}$ and high-$M_{2}$. However, the threshold choice is a way to parametrizes the difficulty of the classification task. In particular, for low $M_{2}$ threshold the SRE-based classification becomes a physics-motivated problem where the goal is to distinguish between stabilizer and magic states on a real quantum hardware. In this case, the threshold will respresent an indication of the quantum noise impact on the $M_{2}$ value of the stabilizer states.
4.3 Data for Stabilizer Rényi Entropy Estimation
The procedure used to generate the dataset for the regression formulation of the SRE estimation problem follows the approach described in [lipardi2025study], and reported in the following. In addition, we extend the dataset to noisy quantum simulations to investigate the effect of quantum noise on the performance of our model.
The dataset comprises two classes of quantum circuits, both defined for systems with a number of qubits $n$ ranging from $2$ to $6$. The first class includes $250000$ random quantum circuits (unstructured), $50000$ for each value of $n$, and is hereafter referred to as the RQC dataset. The second class includes $25000$ quantum circuits based on the Trotterized dynamics of the one-dimensional transverse-field Ising model (structured), $5000$ for each value of $n$, hereafter referred to as the TIM dataset. All circuits are labeled with their $M_{2}$ value, computed from Equation 4. Each circuit is associated with two labels: the first corresponds to the $M_{2}$ obtained from noiseless simulations using a statevector simulator, and the second to the $M_{2}$ obtained from noisy simulations emulating the FakeOslo backend of IBM Quantum. This dual labeling enables a systematic analysis of the impact of quantum noise on the GNN model performance.
@@ 4.3 Data for Stabilizer Rényi Entropy Estimation
RQC Noisy 100000 262\!-\!6 1801\!-\!80 Qubits and Gate count
@@
TIM Noisy 15000 262\!-\!6 151\!-\!5 Qubits and Trotter steps
These datasets are designed to be representative of typical circuits used in applications for NISQ devices [preskill2018quantum], covering a wide range of $M_{2}$ values. The variety of quantum circuit structures included in the dataset is crucial for studying the generality and robustness of our approach. On the one hand, random quantum circuits provide a general and unstructured benchmark, as they are not linked to any specific problem domain. On the other hand, the transverse-field Ising model serves as a well-established benchmark in quantum machine learning, owing to its fundamental role in condensed matter physics, statistical mechanics, and quantum information. Moreover, the behavior of the SRE in the transverse-field Ising spin chain has been extensively investigated [oliviero2022magic].
Table 2 provides a schematic overview of the datasets used for the SRE estimation problem. Figure 3 illustrates the $M_{2}$ value distribution under noiseless simulations for the RQC and TIM datasets. Specifically, Figures 3(a) and 3(b) show the $M_{2}$ distributions of quantum circuits grouped by qubit count, with their average values indicated by dashed vertical lines, for the RQC and TIM datasets, respectively. Figures 3(c) and 3(d) presents the average $M_{2}$ values, together with their standard deviations (shaded regions), for quantum circuits grouped by qubit number as a function of circuit depth. On the other hand, Figure 4 shows the corresponding distributions obtained from noisy simulations in comparison to the noiseless case. Figures 4(a) and 4(b) demonstrate that noisy quantum simulations have a stronger impact on the TIM dataset than on the RQC dataset: on average, the $M_{2}$ values in the TIM dataset shift more towards higher values, and the overall distributions become more stretched.
(a) RQC - SRE Distribution
(b) TIM - SRE Distribution
(c) RQC - $M_{2}$ vs Gate count
(d) TIM - $M_{2}$ vs Trotter steps
@@
where $Z_{i}$ and $X_{i}$ are Pauli operators acting on qubit $i$, $J$ denotes the interaction strength, and $h$ the transverse-field strength. The time-evolution operator $U(t)=e^{-iHt}$ is simulated using a first-order Trotter–Suzuki decomposition [nielsen2010quantum]. Each Trotter step is constructed as a sequence of quantum gates: the two-qubit interaction $e^{-i\theta Z_{i}Z_{i+1}}$, with $\theta=J\Delta t$, is realized by the standard $\texttt{CNOT}_{i,i+1}$$\texttt{RZ}(2\theta)_{i+1}$$\texttt{CNOT}_{i,i+1}$ construction, while the transverse-field term $e^{-i\phi X_{i}}$, with $\phi=h\Delta t$, is implemented as a single-qubit rotation $\texttt{RX}(2\phi)_{i}$. The circuits span various values of the angle parameters $\theta$, $\phi$, with the number of Trotter steps ranging in $T=[1,5]$.
@@
Figure 4: SRE value distributions of quantum circuits grouped by qubit number in the RQC and TIM noisy datasets are shown in Figure 4(a) and 4(b), respectively. The $M_{2}$ values are computed under simulated noise using the IBM FakeOslo backend.
5 Graph Neural Networks for Nonstabilizerness Estimation
This section describes the proposed Graph Neural Network approach for nonstabilizerness estimation under both classification and regression problem formulations. All problem formulations share the same graph-based quantum circuit representation and GNN architecture. Figure 5 provides an overview of the whole pipeline. The proposed graph-based representation consists of two main components. The first, in yellow, is the directed acyclic graph representation of the circuit, which is encoded as an adjacency matrix together with node embeddings that represent quantum gates. The second component, in green, is the global feature set encoding the total gate count. These two components are processed independently, the first is passed through a sequence of Transformer convolutional layers, while the second through a fully connected neural network. The resulting latent representations are concatenated and passed through a regressor. Section 5.1 and Section 5.2 describe the details of the encoding technique employed and the GNN architecture, respectively.
Figure 5: Overview of the GNN approach to nonstabilizerness estimation. Box A summarizes the three different problem formulations, box B illustrates the process to translate a quantum circuit in our graph-based representation, and box C illustrates the GNN architecture.
5.1 Quantum Circuit Representation
Quantum circuits can be naturally represented by directed acyclic graphs, where each node is a gate and each edge is a qubit acted upon by a gate [liao2024machine, aktar2024graph]. The graph passed in input to the GNN is encoded by two matrices, the adjacency matrix $A$ representing the graph structure, and the node feature matrix $X$, which contains the node embeddings. Since we deal with directed graphs, the adjacency matrix is not symmetric. Each node has an associated node embedding $x\in\mathbb{R}^{d}$. The size of the node embedding is $d=d_{g}+d_{q}+d_{h}$, as it stores the information regarding the gate type in $d_{g}$ components, the target qubit(s) upon which they are applied in $d_{q}$ components, and the specifics of the quantum backend in $d_{h}$ components,. The gate type and qubit indices are encoded using one-hot encoding, while the quantum backend vector is intrinsically real-valued. The one-hot encoding for the gate type requires $d_{g}=7$ bits, one for each gate type: input gate (qubit initialization), output gate (measurement), CNOT, H, RX, RY, and RZ. The one-hot encoding for the target qubit requires a number of bits equal to the number of qubits $d_{q}$ in the largest quantum circuit of the dataset. The hardware-dependent noise specifics of the gate applied require a real-valued vector of $d_{h}=7$ seven components, encoding the relaxation time $T^{q}_{1}$, the dephasing time $T^{q}_{2}$ and readout error $r_{q}$ on the qubit(s) $q$ on which the gate is applied, and the gate error $g_{\epsilon}$. Thus, in the stabilizer state classification and the SRE-based classification, where the datasets include quantum circuits with up to $25$ qubits the dimension of the node embedding is $d=7+25+0=32$. In the noiseless SRE estimation task, where circuit have up to $6$ qubits the embedding dimension is $d=7+6=13$. In the noisy SRE estimation, we concatenate to the base node embedding of the noiseless case the $7$-dimensional vector encoding the specifics of the IBM Fake Oslo backend, resulting in $d=13+7=20$. Figure 5 provides a comprehensive overview on the whole pipeline, including the architecture of the model.
The global feature set provided as input to the GNN model is a vector that encodes the gate count of quantum circuits. This representation has proven effective for supervised learning approaches to SRE estimation [lipardi2025study]. It consists of a vector of length $152$, where unparameterized gates are directly counted, whereas parameterized gates are discretized by dividing the interval $0,2\pi[$ of the angle parameter into fifty equally spaced bins. Figure 6 illustrates both the global feature set and the node embeddings of the graph-based circuit representation.
5.2 Graph Neural Network Architecture
The architecture of the Graph Neural Network employed in this work is inspired by previous works, where GNNs models are trained to predict the expressibility of quantum circuits [aktar2024graph] and to predict noise impact on circuit fidelity [wang2022torchquantum].
The model is divided into two main independent components. The first component processes the graph representation, including the adjacency matrix $A$ and the node embeddings $X$, which are initialized as vectors encoding the gate type and target qubit(s), as described in the previous section. These are processed by three Transformer Convolutional (TC) layers that aggregate local neighborhood information within the graph. A ReLU activation function is applied between each pair of consecutive TC layers to mitigate potential numerical issues arising from negative values [aktar2024graph]. Finally, the resulting feature vectors, obtained after applying the sequence of TC and ReLU layers, are aggregated using global mean pooling. The second component is a fully connected artificial neural network that takes as input the global feature set, the vector encoding the gate counts of entire quantum circuits, as defined in the previous section. Finally, the output vectors from the two components are concatenated and passed through a regressor consisting of three fully connected layers interleaved with two ReLU activation layers.
However, there are key differences between the GNN implementations for the classification and regression tasks. In the stabilizer state classification setting, we use the binary cross-entropy loss function, and the output of the GNN is passed through a sigmoid function with a threshold of $0.5$ to assign the class label of each quantum circuit. In contrast, for the SRE estimation problem, we employ the Huber loss function, and the final output of the GNN directly represents the predicted $M_{2}$ value of the input quantum circuit. Additionally, the circuit representations differ slightly across the various experimental configurations, as described in the previous section. Note that we employed a grid search to tune some hyperparameters of the GNN, such as the number of the TC layers, the hidden dimension of the fully connected neural networks, and the learning rate.
6 Experimental Results
This section reports the experimental setup and results conducted to assess the proposed GNN approach on three different problem formulations of increasing difficulty. Section 6.1 focuses on the stabilizer state classification task, Section 6.2 on the more challenging SRE-based classification task, and Section 6.3 on the SRE estimation task. Each section outlines the training procedure and performance evaluation by assessing the generalization capability of the proposed GNN in predicting $M_{2}$ across different scenarios. when tested on quantum circuits with a higher number of qubits, a larger total gate count, and greater entanglement levels than those used for training.
Section 6.4 investigates the sources of improvement achieved by our GNN relative to prior SVR-based approaches [lipardi2025study]. To this end, we evaluate the model’s performance following the ablation [cohen1988evaluation, newell1975tutorial] of the graph-based representation and the Transformer layers that process it, resulting in a simplified fully connected neural network trained only on the global feature set. Finally, Section 6.5 extends the SRE estimation experiments to a noisy setting, considering transpiled quantum circuits labeled by simulating the noise of an IBM quantum backend.
6.1 Stabilizer State Classification
The experiments for training and testing our GNN model on the stabilizer state classification problem can be divided into two main parts. The first part focuses on quantum circuits with a fixed number of qubits, set to $18$, to allow a direct comparison with the previous approach [mello2025retrieving]. The second part extends the experiments to a broader dataset of quantum circuits with the number of qubits ranging from $2$ to $25$.
In the first part of the experiments, GNN is trained on $10000$ out of the $15000$ circuits in PS 18, where both the training and test set are balanced between the two labels, stabilizer and magic states. Our GNN achieves a classification accuracy of $99.98\%$ on the training set and $99.72\%$ on the test set. When evaluated on the Clifford-evolved circuits of the dataset CS 18, the GNN accuracy keeps stable for the classification of stabilizer states, while it decreases as a function of the Clifford depth for magic states, see the blue lines in Figure 7. This phenomenon is also experienced by the CNN model. However, the classification accuracy obtained by the CNN decreases when evaluated on Clifford-evolved quantum states with Clifford depth ranging from 0 to $8$. For magic states, it decreases from $100\%$ to approximately $90\%$, and for stabilizer states, from $100\%$ to $80\%$. In contrast, our GNN manages to maintain stable the perfect classification accuracy for stabilizer states, and for magic states the accuracy only decreases from almost $100\%$ to $99\%$ at depth $8$. Furthermore, we extended the analysis up to a Clifford depth of $25$, where the accuracy obtained by the GNN on magic states decreases to $97\%$, while the accuracy on the stabilizer states is still stable to $100\%$. We refer the reader to Figure 7.
(a) Trained on PS 18
(b) Trained on PS 2-10
@@ 6.1 Stabilizer State Classification
6.2 SRE-based Classification
(a) Generalization across different scenarios
(b) Performance dependence on $M_{2}$
6.3 Stabilizer Rényi Entropy Estimation
The experiments for training and testing our GNN model for SRE estimation are aligned with the setting proposed in [lipardi2025study]. The experiments can be divided into two main parts. The first part is designed to evaluate the generalization capabilities of GNN in terms of number of qubits, while the second in terms of circuit depth.
To assess the generalization capabilities of our GNN approach, we test its performance on out-of-distribution instances. In particular, we evaluate how the GNN generalizes with respect to the number of qubits and the gate counts. We refer to these experiments as extrapolation in qubit number and in gate count, respectively. We train our GNN independently on the two quantum circuits datasets, RQC and TIM. In all experiments we split training and test set in 8080-2020. Note that, to ensure statistical robustness, each experiment is repeated ten times with different random seeds, and the reported results correspond to the average results over these ten independent runs.
Figure 12 summarizes the experiments designed to test the extrapolation in the qubit number. Here, the models are trained on the subset of quantum circuits with qubit numbers from $2$ to $5$, and then evaluated on $6$-qubit circuits. The difference between the MSE values obtained by the SVR on the training and test set suggests the presence of overfitting. This effect is observed in both datasets. In the RQC dataset, the MSE increases from $0.175$ on the training set to $0.202$ on the test set, while in the TIM dataset, the MSE increases from $0.22$ to $0.25$. As a result, the generalization performances are limited as the MSE increases of approximately of a factor $3.5$ in the RQC dataset and $4.5$ in the TIM dataset. In contrast, the GNN model significantly improves upon the results obtained with SVR, particularly in terms of Mean Squared Error (MSE) on the extrapolation set. Specifically, GNN reduces the extrapolation MSE upon SVR by the $79\%$ on RQC, and by $26\%$ on TIM. The highest improvement is achieved on the RQC dataset, where the GNN also reduces the gap between the MSE values obtained on the training and test set.
Figure 13 summarizes the experiments designed to test the extrapolation in gate counts. Here, the models are trained on the subset of circuits with gate counts restricted to the range $[0,79]$ for the RQC dataset and evaluated on circuits with gate counts ranging in $[80,99]$. This approach is similarly applied to the TIM dataset, where the model is trained on circuits with trotter steps from $1$ to $4$ and tested on trotter step $5$. The difference between the MSE values obtained by the SVR on the training and test sets indicates a slight overfitting effect. This behavior is more evident on the RQC dataset, where the MSE increases from $0.189$ in the training set to $0.219$ in the test set. In contrast, for the TIM dataset, the MSE increases only from $0.26$ to $0.28$, resulting in a slightly better generalization performance compared to the RQC dataset. In this setting, our GNN demonstrates strong generalization performance. The gap between training and test MSE observed in the SVR model for the RQC dataset is reduced, and concurrently the extrapolation MSE improves by $95\%$, from $0.532$ to $0.025$. Moreover, the GNN achieves an extrapolation MSE of $0.028$, which improves the SVR model by $56\%$. Table 3 summarizes the results obtained in terms of MSE on the extrapolation set compared to the Support Vector Regressor proposed in [lipardi2025study]. In addition, it includes the results obtained by training a Neural Network as described in Section 6.4.
@@ 6.3 Stabilizer Rényi Entropy Estimation
It is important to focus on the extrapolation MSE obtained on the RQC dataset when extrapolating on gate counts, as it is unexpectedly lower than the training MSE. However, Figure 3(c) shows that the SRE distribution of the quantum circuits in RQC becomes narrower, with the standard deviation decreasing as the gate count increases. Since $M_{2}$ is bounded above by $SRE_{max}(n)=\ln(\frac{2^{n}+1}{2})$ [leone2022stabilizer], for a quantum system of $n$ qubits, higher gate counts that include non-Clifford gates drive the system towards a more stable regime with smaller variance. To further investigate this effect, we perform an additional experiment in which we train our GNN on quantum circuits with the largest gate counts, namely from $20$ to $99$, and extrapolate on the shallower circuits with gate counts from $1$ to $19$. In this setup, we require the model to extrapolate on a subset of circuits with a wider interval of $M_{2}$, higher standard deviation, and less separation. As a result, extrapolating in the new setup should be more difficult. Figure 14 presents a direct comparison between two extrapolation experiments. The bars on the left, relate to the GNN extrapolation on higher gate counts. On the right, there are the bars related to the extrapolation to unseen circuits with lower gate counts. As expected, the extrapolation MSE in the latter setting is higher than in the former, and also exceeds the training MSE.
Table 3: Relative performance of the GNN and $N$ models compared to the SVR baseline [lipardi2025study], in terms of the extrapolation MSE.
Extrapolation on Qubit Number
Dataset GNN NN
RQC -78.8% +4.3%
TIM -25.6% -55.6%
Extrapolation on Gate Counts
Dataset GNN NN
RQC -95.3% -1.9%
TIM -55.6% -31.6%

### v2

@@ A.1 Dataset for Stabilizer State Classification
The presence of CS circuits in the dataset allows to investigate the generalization capabilities of the models to unseen entangled states, while keeping the same labels and $M_{2}$ values, in line with the procedure used in [36]. The Entangled States (ES) subset, on the other hand, enables a further step in this analysis by testing model performance on circuits characterized by substantially different structural patterns and, in general, a higher degree of entanglement. Figure 10 shows the heterogeneous quantum circuit dataset generated with our procedure, spanning a broad range of qubit number, circuit depth, circuit structure, entanglement and $M_{2}$. Figure 10(a) illustrates the diversity of quantum circuits across the three different types in terms of gate counts and $M_{2}$. It is important to note that our procedure generates also quantum circuits with significantly low $M_{2}$ values. For instance, in the case of $18$-qubit circuits, the minimum $M_{2}$ value obtained in our dataset is $0.01$, compared to $1.80$ reported in [36]. Moreover, in the dataset PS 2-25, we observe a minimum $M_{2}$ value of $0.006$, see Figure 10(b). This is relevant, as the performance of machine learning models on the stabilizer state classification problem is typically correlated with the density of $M_{2}$ in the number of qubits $n$: $m_{2}=\frac{M_{2}}{n}$. Models tend to achieve higher accuracy for states with higher $m_{2}$, because of the increased separability of the magic states from the stabilizer states. This effect is also shown in the experimental results discussed in [36] and further confirmed by our own results obtained using graph neural networks (GNNs).
@@
A.2 Dataset for SRE-based Classification
Here, we provide a further analysis of the dataset generated for the SRE-based classification problem, which we recall it results from a simple procedure applied to the dataset generated for the stabilizer state classification problem. Specifically, the stabilizer states (labeled as 0) are removed from the dataset, and among the remaining circuits (originally labeled as $1$), those with $M_{2}$ below a given threshold are relabeled as 0. The violin plots in Figure 11 illustrates the distribution of $M_{2}$ values across the product-state datasets. The horizontal lines in the plot represent the median $M_{2}$ value for each dataset shown along the x-axis. As expected, quantum circuits with qubit numbers between $2$ and $10$ exhibit a narrower range of $M_{2}$ values compared to circuits with qubit numbers between $11$ and $25$.
@@ A.2 Dataset for SRE-based Classification
We note that by choosing the median as the threshold $M_{2}$ value in our dataset we ensure that is balanced between the two labels: low-$M_{2}$ and high-$M_{2}$. However, the threshold choice is a way to parametrizes the difficulty of the classification task. In particular, for low $M_{2}$ threshold the SRE-based classification becomes a physics-motivated problem where the goal is to distinguish between stabilizer and magic states on a real quantum hardware. In this case, the threshold will represent an indication of the quantum noise impact on the $M_{2}$ value of the stabilizer states.
A.3 Dataset for Stabilizer Rényi Entropy Estimation
In this section we further analyze the dataset generated for the SRE estimation problem, introduced in Section 4.3. Figures 12(a) and 12(b) presents the average $M_{2}$ values, together with their standard deviations (shaded regions), for quantum circuits grouped by qubit number as a function of circuit depth. On the other hand, Figure 13(b) shows the corresponding distributions obtained from noisy simulations in comparison to the noiseless case. Figures 13(a) and 13(b) demonstrate that noisy quantum simulations have a stronger impact on the TIM dataset than on the RQC dataset: on average, the $M_{2}$ values in the TIM dataset shift more towards higher values, and the overall distributions become more stretched.
Table 3 provides a schematic overview of the datasets used for the SRE estimation problem. Figure 12 illustrates the $M_{2}$ value distribution under noiseless simulations for the RQC and TIM datasets.
(a) RQC
(b) TIM
@@ A.3 Dataset for Stabilizer Rényi Entropy Estimation
RQC Noisy 100000 262\!-\!6 1601\!-\!60 Qubits and Gate count
@@
TIM Noisy 15000 262\!-\!6 131\!-\!3 Qubits and Trotter steps
@@
where $Z_{i}$ and $X_{i}$ are Pauli operators acting on qubit $i$, $J$ denotes the interaction strength, and $h$ the transverse-field strength. The time-evolution operator $U(t)=e^{-iHt}$ is simulated using a first-order Trotter–Suzuki decomposition [38]. Each Trotter step is constructed as a sequence of quantum gates: the two-qubit interaction $e^{-i\theta Z_{i}Z_{i+1}}$, with $\theta=J\Delta t$, is realized by the standard $\texttt{CNOT}_{i,i+1}$$\texttt{RZ}(2\theta)_{i+1}$$\texttt{CNOT}_{i,i+1}$ construction, while the transverse-field term $e^{-i\phi X_{i}}$, with $\phi=h\Delta t$, is implemented as a single-qubit rotation $\texttt{RX}(2\phi)_{i}$. The circuits span various values of the angle parameters $\theta$, $\phi$, with the number of Trotter steps $T$ ranging from one to five.
@@
Figure 13: SRE value distributions of quantum circuits grouped by qubit number in the RQC and TIM noisy datasets are shown in Figure 13(a) and 13(b), respectively. The $\tilde{M_{2}}$
values are computed under simulated noise using the IBM FakeOslo backend.
Appendix B. Complementary Analysis to the Experimental Results
In this section, we provide additional experiments and analysis to Section 6.
B.1 Performance Analysis for the Stabilizer State Classification
@@
B.2 Performance Analysis for the Stabilizer Rényi Entropy Estimation
@@
It is important to focus on the extrapolation MSE obtained by the proposed GNN on the RQC dataset when extrapolating on gate counts, as it is unexpectedly lower than the training MSE. Figure 12(a) shows that the SRE distribution of the quantum circuits in RQC becomes narrower, with the standard deviation decreasing as the gate count increases. Since $M_{2}$ is bounded above by $SRE_{max}(n)=\ln(\frac{2^{n}+1}{2})$ [29], for a quantum system of $n$ qubits, higher gate counts that include non-Clifford gates drive the system towards a more stable regime with smaller variance. To further investigate this effect, we perform an additional experiment in which we train our GNN on quantum circuits with the largest gate counts, namely from $20$ to $99$, and extrapolate on the shallower circuits with gate counts from $1$ to $19$. In this setup, we require the model to extrapolate on a subset of circuits with a wider interval of $M_{2}$, higher standard deviation, and less separation. As a result, extrapolating in the new setup should be more difficult. Figure 15 presents a direct comparison between two extrapolation experiments. The bars on the left, relate to the GNN extrapolation on higher gate counts. On the right, there are the bars related to the extrapolation to unseen circuits with lower gate counts. As expected, the extrapolation MSE in the latter setting is higher than in the former, and also exceeds the training MSE.